Please help me solve the linear system by giving me the cordinates
15x + y =70
3x - 2y =-8

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Please help me solve the linear system by giving me the cordinates
15x + y =70
3x - 2y =-8
=============================
Multiply the 1st eqn by 2, then add.

Answer by stanbon(75887)

Please help me solve the linear system by giving me the cordinates
15x + y =70
3x - 2y =-8
--------
Modify::
30x + 2y = 140
3x - 2y = -8
-----------------
Add and solve for "x":
33x = 132
x = 4
-----
Solve for "y"::
y = -15x + 70
y = -15*4 + 70
y = 10
------------------
Cheers,
Stan H.
------------------